OpenStudy (anonymous):

Find all values of k so that (-1,2), (-10,5), and (-4,k) are the vertices of a right triangle.

OpenStudy (anonymous):

use length formula to find the length of the first two sides and then use pythagorean theorem to finf the length of the last side and substitute the known values and find k

OpenStudy (anonymous):

Thanks @MohamedDada

OpenStudy (anonymous):

i try =D

OpenStudy (anonymous):

what's the length formula ?

OpenStudy (anonymous):

Its sqrt((x1-x2)^2+(y1-y2)^2) anyway do you know slope formula that a better formula to use in this question...

OpenStudy (anonymous):

d = (sqrt(x2 - x1)^2 (y2-y1)^2)

OpenStudy (anonymous):

can you do the first steps for me? I'm seriously confused, geometry is to hard for me :[

OpenStudy (anonymous):

msn?

OpenStudy (anonymous):

i don't have msn

OpenStudy (anonymous):

ok ill just do it here

OpenStudy (anonymous):

okay [:

OpenStudy (anonymous):

\[d = \sqrt{(-10 + 1) (5 - 2)}\]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Err that is not correct it is sqrt(9^2+3^2)=sqrt(90).

OpenStudy (anonymous):

i know -.-

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

which one iisnt rght ?

OpenStudy (anonymous):

i forgot to square

OpenStudy (anonymous):

Ok i'll tell you how to do it THe product of slope's of the lines which are perpendicular are -1 so slopes of the three lines are 1. (k-5)/6 2. 3/-9=-1/3 3. (k-2)/(-3)=(2-k)/3 So you take them in pairs and apply the given conditon that product is -1. I'll do one for you i'll take 1 and 3 and mutiply equating to -1, (k-5)(2-k)/(6*3)=-1, k^2-7k+-8=0 (k-8)(k+1)=0 k=8,k=-1. These are two values now you take the remaining two pairs and do it you'll get two more vaules.

OpenStudy (anonymous):

Do you know slope formula?

OpenStudy (anonymous):

i'm not sure is this all the steps? o.o

OpenStudy (anonymous):

Slope formula slope of a line is (y2-y1)/(x2-x1) So the first slope is(k-5)(-4-(-10)) you get (k-5)/6.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

how did you get the (-8) ??? i did it several times and i get -9 !!!